OpenStudy (anonymous):

Please help!! Consider the function f(x)=(3x^2)-(2x^-1)-(3x^1/3)+2 Determine the equation of the line tangent to y=f(x) at x=-1

OpenStudy (baldymcgee6):

Do you know how to take the derivative?

OpenStudy (anonymous):

Yes!

OpenStudy (baldymcgee6):

Show me!

OpenStudy (anonymous):

I used the power rule and got - f'(x)= (6x)+(2x^-2)-(x^-2/3)

OpenStudy (baldymcgee6):

you should get \[y'=2/x^2-1/x^{2/3}+6 x\]

OpenStudy (baldymcgee6):

yep same thing!

OpenStudy (anonymous):

Okay, so then how do I do the other question? :S

OpenStudy (baldymcgee6):

So with the derivative we found, we can find the SLOPE of the line at x = -1

OpenStudy (baldymcgee6):

f'(-1)= (6(-1))+(2(-1)^-2)-((-1)^-2/3) =???

OpenStudy (baldymcgee6):

tell me what you get for our slope

OpenStudy (anonymous):

-5?

OpenStudy (baldymcgee6):

Make sure you type it into your calculator properly.. I got -9

OpenStudy (anonymous):

I tried a couple time and also asked my friend to try and we both get -5.

OpenStudy (baldymcgee6):

hahah oops. my bad, it is -5

OpenStudy (baldymcgee6):

from here, we know that the slope of the tangent line is -5, therefore y=mx+b y = (-5)x+b y = (-5)(-1)+b Notice we still have two unknowns,is there a way we can get a y value from some where?

OpenStudy (baldymcgee6):

if we use f(x)=(3x^2)-(2x^-1)-(3x^1/3)+2, where x = -1, we will get an output of the y co ordinate.. so, find f(-1) now

OpenStudy (anonymous):

Well I also had to find f(-1) which equals 10. So is that my y?

OpenStudy (baldymcgee6):

yes it sure it, it follows that, y = mx+b 10 = (-5)(-1) + b Solve for b

OpenStudy (anonymous):

Okay, thank you! :)

OpenStudy (baldymcgee6):

so your tangent line is y = -5x + 5 Please leave fan and best response! :)